detremine whether the graphs of each pair of lines are parallel
3x+8=y
2y=6x-3

Answer by ewatrrr(10682)


Hi
the standard slope-intercept form for an equation of a line is y = mx + b
where m is the slope and b the y-intercept.
Solving for y to put lines ito the slope-intercept in order to read their slopes
3x+8=y y = 3x + 8 slope m = 3
2y=6x-3 y = 3x - 3/2 slope m = 3
Lines have identical slopes, therefore, are parallel